OpenStudy (anonymous):

What are the x- and y – intercepts of the following equation: y = 60x + 300 Answers: (0, 300), (-5, 0) (0, 300), (0, -5) (300, 0), (5, 0) (300, 0), (0, 5)

OpenStudy (johnweldon1993):

To find the x-intercept.....replace 'y' in your equation with 0....and solve for 'x' To find the y-intercept....replace the 'x' in your equation with 0...and solve for 'y'

OpenStudy (anonymous):

so x= 60(0) & y=0 @johnweldon1993

OpenStudy (johnweldon1993):

Not quite y-intercept = replace 'x' with 0 so y = 60x + 300 After replacing we have y = 60(0) + 300 simplify y = 0 + 300 or just y = 300

OpenStudy (anonymous):

And x-intercept = replace 'y' with 0 so y = 60x + 300 After replacing... y = 0 + 300 (simplified) or just y = 300

OpenStudy (johnweldon1993):

No you replaced the 'x' with 0 again it should have been y = 60x + 300 replace the 'y' with 0 so 0 = 60x + 300 Now solve for 'x' I'll get you started 60x = -300 x = ...?

OpenStudy (anonymous):

60(0) ???

OpenStudy (anonymous):

OR -300

OpenStudy (johnweldon1993):

No now you just solve for 'x' 60x = -300 divide both sides by 60 x = -5 Here let me do it again You have the equation \[y = 60x + 300\] To find the y-intercept....you replace the 'x' with 0....because at the y-intercept on a graph x DOES = 0 right? so After replacing the 'y' in the equation with 0....we have \[0 = 60x + 300\] Now we just solve for 'x' Subtract 300 from both sides... \[60x = -300\] And divide both sides of the equation by 60 \[x = -5\] So this is your x-intercept To solve for the y-intercept...we replace the 'x' in the equation with 0...and solve for 'y' After replacing we get \[y = 60(0) + 300\] Now we know 60 times 0 is 0 so this is simply \[y = 300\] Does that make any more sense?

OpenStudy (anonymous):

yes thank you so much .... ! @johnweldon1993

OpenStudy (johnweldon1993):

No problem! So what would be your answer choice?

OpenStudy (anonymous):

A

OpenStudy (johnweldon1993):

Good work!

OpenStudy (anonymous):

thnx